Parth (parthkohli):

A very basic tutorial. How to deduce that \(0 \times 2 = 0 \)? 1) The translation for \(a \times b\) when \(b \in \mathbb{N} \) is \(a + a + a \cdots (\text{b times})\) Therefore, \(0 \times 2 = 0 + 0\). Going a little elementary — No apples and no apples are no apples. 2) Zero is considered as the additive identity in Mathematics. The property says: \(x + 0 = x\) Assume that \(x = 0\), so \(0 + 0 = 0\)

Parth (parthkohli):

Source: I posted this in M.Se

Parth (parthkohli):

Refresh the page, I edited the thing.

OpenStudy (anonymous):

It's via definition.

Parth (parthkohli):

This might not be good for M.Se, but it's cool for OS :P

OpenStudy (unklerhaukus):

should this arrow really go both ways?\[x + 0 = x\Longleftrightarrow x \in \mathbb{R}^{+}\]

Parth (parthkohli):

Yeah.

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

Consider \(e+ei.\)

Parth (parthkohli):

Let me edit my tutorial then :O what is the correct notation btw?

OpenStudy (unklerhaukus):

the additive identity only holds for positive real numbers?

Parth (parthkohli):

R+ are the extended real numbers, right?

OpenStudy (anonymous):

\(x+0=x\) holds for all numbers \(x.\) (Exclude set theoretic-numbers; I am uncertain about them.)

Parth (parthkohli):

How do you denote all numbers?

OpenStudy (anonymous):

You don't. As far as I know. I think \(\mathbb{C}\cup \mathcal{T}\) would be the best way?

Parth (parthkohli):

Oh, well. We could just write \(x + 0 = x\). And as I said, this is a basic tutorial.

OpenStudy (anonymous):

\(\mathcal{T}\) for transcendentals.

OpenStudy (unklerhaukus):

\[x + 0 = x\Longleftarrow x \in \mathbb C\] \[\mathbb R \subset\mathbb C\]

Parth (parthkohli):

My tutorial didn't go as well. I closed it.

OpenStudy (unklerhaukus):

just put \[\Longleftarrow\] instead of \[\Longleftrightarrow\]

OpenStudy (anonymous):

Rhaukus, you think that transcendentals aren't the same under the additive identity...?

Parth (parthkohli):

I blame @Limitless for telling me that "if and only if" is "\iff or \Longleftrightarrow" :P

OpenStudy (unklerhaukus):

i dont think iff is right

OpenStudy (anonymous):

I mean, it's true that \(x \in \mathbb{C} \Rightarrow x+0=x\), but it's a tad misleading.

OpenStudy (anonymous):

I would think that the most general statement would be \(x \in \mathbb{C}\cup \mathcal{T} \Rightarrow x+0=x.\)

Parth (parthkohli):

Well. Leave it. This tutorial is closed. ;)

OpenStudy (unklerhaukus):

i dont know anything about the transcendentals

OpenStudy (anonymous):

@ParthKohli, no offense intended, but I'm genuinely interested in what Rhaukus's thoughts are.

Parth (parthkohli):

Yeah. That is why I thought that this tutorial failed :)

OpenStudy (unklerhaukus):

this tutorial was a learning experience